A man is 5 years older than his wife who is ten times older than her daughter if her daughter attains the age of 8 in 3 years.what is the age of man?
1
Expert's answer
2012-07-13T07:36:10-0400
Let's denote current age of a daughter by x. Then x+3=8 => x=5.
wife is ten times older than her daughter so the wife is 5*10=50 years old. Then age of a man is 50+5=55